### v1

# CAPE: A CLIP-Aware Pointing Ensemble of Complementary Heatmap Cues for Embodied Reference Understanding

Abstract
We address the problem of Embodied Reference Understanding, which involves predicting the object that a person in the scene is referring to through both pointing gesture and language. Accurately identifying the referent requires multimodal understanding: integrating textual instructions, visual pointing, and scene context. However, existing methods often struggle to effectively leverage visual clues for disambiguation. We also observe that, while the referent is often aligned with the head-to-fingertip line, it occasionally aligns more closely with the wrist-to-fingertip line. Therefore, relying on a single line assumption can be overly simplistic and may lead to suboptimal performance. To address this, we propose a dual-model framework, where one model learns from the head-to-fingertip direction and the other from the wrist-to-fingertip direction. We further introduce a Gaussian ray heatmap representation of these lines and use them as input to provide a strong supervisory signal that encourages the model to better attend to pointing cues. To combine the strengths of both models, we present the CLIP-Aware Pointing Ensemble module, which performs a hybrid ensemble based on CLIP features. Additionally, we propose an object center prediction head as an auxiliary task to further enhance referent localization. We validate our approach through extensive experiments and analysis on the benchmark YouRefIt dataset, achieving an improvement of approximately 4 mAP at the 0.25 IoU threshold.
1 Introduction
Embodied Reference Understanding (ERU) [11] is the task of identifying a specific object in a visual scene based on language instructions and pointing cues in the image. It plays a key role in real-world applications like human-robot interaction, assistive robotics, and augmented reality, where systems must understand which object a person refers to.
While grounding models [37] have made significant progress in detecting objects mentioned in natural language, they often fall short in ERU task, particularly in ambiguous scenes. When multiple instances of the same object type are present, these models tend to detect all matching candidates without the ability to disambiguate and identify the specific target intended by the user. Moreover, when the textual instruction itself is vague or ambiguous, grounding models struggle even further, often failing to identify the correct target object or producing no confident prediction at all due to the inherent ambiguity. This limitation highlights the need for additional disambiguation cues, embodied gesture signals, that can help resolve referential ambiguity and enable accurate identification of the intended object. Therefore, both embodied gesture signals and language reference are crucial to identify the referent. In ERU, two main challenges arise. (1) identifying candidate objects from the text, as in standard grounding tasks, and (2) interpreting the pointing cue from visual input to guide the final prediction. The latter requires understanding human pose, inferring pointing direction, and handling visual complexities like perspective, occlusion, and depth.
In [11], first time ERU task is defined and a multimodal YouRefIt dataset is proposed. They also proposed a method that combines textual and visual inputs along with PAF [7] and predicted saliency maps [33] to help localize the human and narrow down the potential referent. While these inputs lead to improved performance, the model still exhibits suboptimal accuracy in many cases, particularly due to its inability to properly follow the pointing instructions. Building on this, the Touch-Line Transformer [35] introduces a novel approach by predicting a Virtual Touch-Line and applying losses that encourage the model to focus on the head-to-fingertip direction, which they find often aligns closely with the intended referent [24, 44]. Although this method achieves significant gains, it still faces challenges. The model occasionally overlooks the pointing line, as it is not explicitly provided as input. Instead, it predicts head, fingertips, and arm position to encourage attention in that direction. More importantly, the head-to-fingertip direction does not always align with the target. In some scenarios, the accurate pointing direction follows a line from the hand alone (e.g., wrist-to-fingertip), rather than from the head (Fig. 1). These observations suggest that while head-to-fingertip direction offers valuable guidance, it is not sufficient on its own to fully resolve the complexity of the task.
To tackle these challenges, we propose a dual-model framework that leverages complementary pointing cues. Both models take as input a referring expression and an image with a pointing gesture, along with an additional heatmap that encodes the pointing direction. One model utilizes a head-to-fingertip heatmap, while the other uses a wrist-to-fingertip heatmap, capturing different aspects of the pointing behavior. Their predictions are integrated via a CLIP-Aware Pointing Ensemble (CAPE) module, which effectively combines the strengths of both models. Additionally, we introduce an auxiliary objective to predict the object center, providing an extra supervisory signal that further enhances grounding accuracy. Our contributions are: (1) We propose using heatmap as extra modality to guide our end-to-end reference object detection model to focus on pointing information more. (2) We propose two parallel complementary models to focus on head-to-fingertip and wrist-to-fingertip pointing lines, to individually alleviate inherent challenges of the task. (3) We present CAPE module to effectively ensemble two complementary models. (4) We introduce object center prediction head to enhance prediction accuracy by focusing on finding the object without considering the precise bounding box.
2 Related Work
Referring Expression Comprehension (REC)
In REC [23, 31, 42, 46, 47, 72], the goal is to identify a specific region in an image based on a given referring expression. In contrast to traditional object detection, REC interprets free-form text and can locate objects from any category, including previously unseen ones. Similarly, OV-DETR [73] integrates image and text embeddings from a CLIP model as queries within the DETR [9] to generate category-specific bounding boxes. ViLD [19] distills knowledge from a CLIP teacher model into an R-CNN-like detector [18], enabling region embeddings to capture semantic information from language. GLIP [17] formulates object detection as a grounding task, using additional grounding datasets to align visual regions with textual phrases. DetCLIP [71] enriches its knowledge base using generated pseudo labels. While YOLO-World [12] extends traditional YOLO object detection [50] to an open-world setting, GroundingDINO [37] aligns visual and textual features to detect arbitrary objects described by free-form text. Recently, LMMs [5, 1, 3] get attention for their superior performance in visual grounding [10].
Nonverbal Communication for Referent
Some studies [55, 56, 57, 67, 6, 66, 25, 45, 61, 29, 53, 64, 63, 20, 60, 68, 58, 62, 59, 21, 51] use gaze target detection as a nonverbal cue, however, gaze alone is unreliable due to distractions and the absence of a clear pointing moment. It is typically localized frame by frame and often used to support conversation [4, 8]. In [43], a multimodal exophora resolution method is proposed to disambiguate demonstrative expressions like “that one” by integrating object categories, pointing gestures, and prior environmental knowledge. [14] introduced an interactive robot dialogue system that uses multimodal interaction and pointing line estimation to accurately identify referent with an iterative correction process. Their follow-up work [15] extended this by recognizing unseen objects using a Region Proposal Network and VL-T5 multimodal network [13], moving beyond general object classes from benchmarks like MSCOCO [36] to attribute-based object recognition. On the other hand, several datasets have been collected to incorporate nonverbal cues, but most remain in simulated domains due to accessibility and controllability [26, 27, 2, 28].
Embodied Reference Understanding (ERU)
ERU is a recent advancement of REC that considers the subject’s position while pointing at a reference object. Chen et al. [11] introduced the task and benchmarks using both verbal (text) and nonverbal (pointing gesture) cues. Their model leverages predicted saliency maps and PAF [7] as gestural features to better perceive pointing direction and detect referents. To address ambiguities caused by camera perspective, [52] mapped scenes to 3D coordinates using depth estimation and subject position for spatial attention, but this only yielded modest improvements. Li et al. [35] further improved ERU using an MDETR-based [30] model, introducing a “virtual-touch-line” from eye to fingertip and predicting its vector alongside bounding boxes. In this work, we also use a transformer-based multimodal object detector similar to MDETR [30]. However, instead of predicting a single line vector, we extend the visual-touch-line concept to a pointing heatmap representing the focus of attention. To capture different pointing variations, we use two different heatmaps processed in parallel networks. Some recent works [39, 40] have extended ERU task to 3D embodied settings. ScanERU [39] introduced the first 3D-ERU benchmark by inserting human avatars into existing 3D datasets, but relied on manual placement. In contrast, Ges3ViG [40] improves realism and scalability by automating avatar insertion, generating gesture-aware instructions, and incorporating human localization into the grounding pipeline.
3 Methodology
Problem Definition In ERU, given an RGB image and a text input , we predict a bounding box corresponding to an object referenced by the text and indicated by a pointing gesture in the image. We propose using an additional pointing heatmap to supervise the model better focus on the visual pointing reference.
3.1 Model Architecture
Fig. 2 illustrates our architecture, which consists of a large-scale text encoder, image encoder, and heatmap encoder. The outputs of these encoders are concatenated and fed into a transformer [65] encoder block. Subsequently, a transformer decoder processes the encoder output to generate predictions through task-specific heads. In this work, we introduce two networks that share the same architecture but are trained independently. The first network () receives a heatmap generated from the wrist-to-fingertip line, while the second () uses a heatmap derived from the head-to-fingertip line. Finally, a CLIP-Aware Pointing Ensemble (CAPE) module combines the outputs of these two complementary models to enhance overall performance.
Encoders
We use a pretrained ROBERTA [38] which is a robustly optimized version of BERT [16] as the text encoder to obtain textual embeddings from the input text. To obtain image embeddings, , we utilize a pretrained ResNet-101 [22]. Since the pointing heatmap () is a sparse representation, we choose a lightweight encoder, ResNet-18 [22], to embed it: This choice not only improves the computational efficiency of the network but also helps prevent redundant information in the final concatenated feature representation, which could otherwise harm the model’s learning capacity.
Transformer encoder
After obtaining the embeddings from three encoders, we first apply a convolution to each embedding individually to project their channel dimensions to . Next, we flatten the spatial dimensions of the embedding into a single dimension, converting them into sequences of tokens. For example, for the heatmap embedding: . We then concatenate the three embeddings at the sequence level to form the final representation, which is fed into the transformer encoder. It processes this concatenated input to learn multimodal representations.
Transformer decoder
We feed the multimodal representation output from the transformer encoder into the transformer decoder. Additionally, we provide a set of learnable object queries and gestural keypoint queries. The transformer decoder is responsible for generating object output embeddings and gestural output embeddings.
Prediction head
The object and gestural output embeddings produced by the transformer decoder serve as inputs to our prediction heads. These heads are responsible for predicting candidate bounding boxes for the referent, gestural keypoints (pointing direction prediction head in Fig. 2), which predicts eye, fingertip, wrist coordinates as well as arm class, and center points of the referent. Finally, we select one bounding box prediction, one center point and a pair of gestural keypoints with the highest confidence scores as the final prediction. Please note that we use multi-layer perceptrons (MLPs) as the prediction heads.
3.2 Pointing Heatmap Modality Learning
Incorporating a heatmap that represents the pointing direction provides a valuable spatial priors for the network, especially considering the fact that the existing models in this field tend to give less attention to the visual pointing. This explicit encoding pointing direction helps localize the target region by guiding the model’s attention toward direction and areas that are more likely to be referenced. This is particularly beneficial in real-world, cluttered scenes where gesture interpretation is inherently ambiguous.
How should we create heatmap?
The VTL [35] demonstrates superior performance in most cases, as it effectively captures the intended pointing direction by leveraging both the fingertip position and the head orientation. Human pointing gestures are typically guided by visual attention: we often align our finger with the target object based on our eye-level perspective [35]. Consequently, the head–to-fingertip line serves as a strong proxy for the true pointing intent, integrating both gaze and gesture cues. However, the VTL is not universally reliable. In certain scenarios, the performer may be looking elsewhere (for instance, toward a robot or camera) rather than at the intended object. Additionally, in cases involving extreme viewing angles or occluded head poses, the head orientation may become ambiguous or misleading. In such situations, the local orientation of the hand, as approximated by the wrist–to-fingertip line, can offer complementary directional cues that help disambiguate the pointing gesture. We hypothesize that heatmaps derived from the head-to-fingertip and wrist-to-fingertip lines provide complementary information. Therefore, we generate heatmaps that incorporate both information. For the head-to-fingertip heatmap (), we use the eye coordinates as the starting point and draw a line toward the edge of the image, passing through the fingertip. Similarly, for the wrist-to-fingertip heatmap (), we start from the wrist and draw a line to the edge of the image through the fingertip point. To generate these heatmaps, we apply a Gaussian Ray Heatmap (See Appendix B).
How should we integrate these heatmaps?
The next important question is how to integrate these heatmaps effectively. We explore two different approaches. In the first approach, we merge the two heatmaps () and feed the combined input into a single network to examine whether the model can implicitly learn to utilize both sources of information and determine the most appropriate direction based on the human pose. In the second approach, we train two identical networks independently and ensemble them: one using the head-to-finger heatmap () and the other using the wrist-to-fingertip heatmap (). We empirically find that the second approach yields better performance. As expected, it allows us to train two complementary networks. Notably, the network trained with the head-to-fingertip heatmap () demonstrates significant improvements. In contrast, the first approach results in only subtle enhancement, and the model fails to effectively leverage both sources of information to determine the pointing direction (see Tab. 3 Setup G). Its behavior closely resembles the head-to-fingertip-only scenario. However, due to the presence of additional wrist-to-fingertip line in the merged heatmap, the model experiences contradictory information, ultimately leading to reduced accuracy than head-to-fingertip heatmap-only setup.
Where should we inject these heatmaps?
We empirically find that the most effective way to leverage heatmaps is to use a separate encoder to extract heatmap embeddings, which are then concatenated with the output of the image encoder. It enables us to preserve the pretrained image encoder and allows the dedicated heatmap encoder to extract rich, modality-specific information.
3.3 CLIP-Aware Pointing Ensemble (CAPE)
We introduce an inference time ensemble module to increase the performance by benefiting from two complementary models, and . CAPE is an adaptive scoring method that combines model confidence and CLIP-based similarity in a size-aware manner effectively. (1) For each prediction (top-2 predictions from both models), we compute a normalized CLIP score [48]. We sum pointing models’ confidece scores with the CLIP scores to obtain the fused score (CLIP Fusion in Tab. 5). (2) We calculate CLIP scores for the top-1 predictions of both models. If the confidence scores of the second-highest predictions exceed a threshold, we also compute their CLIP scores. Finally, we select the prediction with the highest CLIP score from among these candidate boxes (CLIP-Only Top-2 + Threshold=0.95 in Tab. 5). CAPE applies strategy (1) when the referent is a small object (defined as occupying less than 0.48% of the image area, following [11]) to avoid relying solely on the CLIP prediction, as CLIP becomes less reliable for smaller objects. For all other objects, strategy (2) is used (see Sec. C.5 for details). This hybrid approach leverages the strengths of both signals while adapting to object scale. All thresholds and selection rules were tuned on the validation set to avoid test-time overfitting.
3.4 Gestural Signal Learning
Referent alignment loss
Following [35], we incorporate a referent alignment loss () to enforce consistency between the predicted referent and the VTL. The core idea is that a correct referent should be geometrically aligned with the pointing direction. To quantify this alignment, we compute the cosine similarity between the eye-to-fingertip vector and the eye-to-object vector, defined as:
where () and () are the eye and the fingertip coordinates. () are the center of the predicted bounding box. We compute the same similarity using GT bounding box to obtain . The referent alignment loss then penalizes deviations between the predicted and GT alignment:
Please note that for both and , we use GT eye and fingertip coordinates to ensure accurate directional representation. We apply this loss during the training of the model , which uses the head-to-fingertip heatmap as reference. For training the model, which utilizes the wrist-to-fingertip heatmap, we replace the eye coordinates with wrist coordinates. The rest remains unchanged, as we similarly aim to maximize the directional correlation.
Object center prediction
In our task, identifying the correct referent and accurately localizing it are two distinct but equally important challenges. To enhance the localization capability of the model, we decouple the prediction of the object center from the rest of the bounding box regression. Specifically, we introduce an additional prediction head implemented as a MLP at the end of the network. This head is dedicated to predicting the () coordinates of the referent object’s center, independently from the standard bounding box regression head. During training, we supervise this center prediction using the L1 loss between the predicted and GT center coordinates:
This explicit supervision encourages the model to focus on spatial alignment and leads to more accurate object localization, especially in scenes with dense or overlapping objects. Moreover, this center prediction serves as an auxiliary task, which improves the representational capacity of the model by encouraging the visual backbone to learn richer geometric and spatial features. We apply this auxiliary head and its corresponding loss to both models. Since this head is used only during training, it introduces no additional computational overhead during inference.
Gesture prediction
In addition to the bounding box and object center predictions, our model incorporates several auxiliary heads to predict task-specific features, including eye (wrist in ) and fingertip coordinates, and arm classification (See prediction head in Sec. 3.1). For each prediction, we apply a standard L1 loss between the predicted and GT data, which encourages spatial attention to task-relevant regions such as the eyes and fingertips, enhancing the model’s understanding of referential gestures. Moreover, we apply cross-entropy loss to classify whether the predicted gestural keypoints are correct.
Total loss
We define our objective as follows:
where is bounding box loss and it is the combination of L1 and GIoU losses as in DETR-like methods [9, 34, 41, 74, 75, 37]. denotes referent alignment loss, indicates object center prediction loss, and is gesture prediction loss. and are soft token loss and contrastive loss respectively to help visual and textural signals alignment as in [30]. We empirically determine the optimal coefficients on the validation set (See Appendix E).
| IoU Threshold for mAP | 0.25 | 0.50 | 0.75 | |||||||||
| Object Sizes | All | S | M | L | All | S | M | L | All | S | M | L |
| Grounding DINO [37] | 57.9 | 38.0 | 60.9 | 74.9 | 54.9 | 35.7 | 59.3 | 69.6 | 42.3 | 22.7 | 45.9 | 58.4 |
| FAOA [69] | 44.5 | 30.6 | 48.6 | 54.1 | 30.4 | 15.8 | 36.5 | 39.3 | 8.5 | 1.4 | 9.6 | 14.4 |
| ReSC [70] | 49.2 | 32.3 | 54.7 | 60.1 | 34.9 | 14.1 | 42.5 | 47.7 | 10.5 | 0.2 | 10.6 | 20.1 |
| YourRefit PAF [11] | 52.6 | 35.9 | 60.5 | 61.4 | 37.6 | 14.6 | 49.1 | 49.1 | 12.7 | 1.0 | 16.5 | 20.5 |
| YourRefit Full [11] | 54.7 | 38.5 | 64.1 | 61.6 | 40.5 | 16.3 | 54.4 | 51.1 | 14.0 | 1.2 | 17.2 | 23.3 |
| REP [52] | 58.8 | 44.7 | 68.9 | 63.2 | 45.7 | 25.4 | 57.7 | 54.3 | 18.8 | 3.8 | 22.2 | 29.9 |
| Touch-Line-EWL [35] | 69.5 | 56.6 | 71.7 | 80.0 | 60.7 | 44.4 | 66.2 | 71.2 | 35.5 | 11.8 | 38.9 | 55.0 |
| Touch-Line-VTL [35] | 71.1 | 55.9 | 75.5 | 81.7 | 63.5 | 47.0 | 70.2 | 73.1 | 39.0 | 13.4 | 45.2 | 57.8 |
| Ours (CAPE) | 75.0 | 63.2 | 80.2 | 81.8 | 65.4 | 49.5 | 74.3 | 72.7 | 35.7 | 13.4 | 40.1 | 53.5 |
4 Experimental Results
Dataset
For both training and testing, we use the image subset of the YouRefIt dataset [11], as the video portion is not publicly available. The dataset contains 2,950 training images and 1,245 test images. Annotations for bounding boxes, as well as eye, fingertip, elbow, and wrist coordinates, are provided by [11, 35].
Evaluation
For fair evaluation and comparison, we follow prior work [11] and report mean Average Precision (mAP) under three Intersection-over-Union (IoU) thresholds: 0.25, 0.50, and 0.75. A prediction is considered correct if its IoU with the GT exceeds the given threshold. Additionally, by following the literature, mAP scores are reported with respect to object size, categorized as Small (S), Medium (M), and Large (L). The object size categories are determined based on the ratio of the object area to the image area, using two thresholds [11]: and . An object is classified as small if its area ratio is below , medium if it falls between and , and large otherwise. In addition to this standard evaluation protocol, we introduce two metrics for assessing referential grounding performance: CLIP score and center coordinate distance (). To compute the CLIP score, we extract features from the text input and the cropped region corresponding to the predicted bounding box using CLIP’s text and image encoders. We then calculate the cosine similarity between these feature vectors to quantify the semantic alignment between the predicted object and the text. For , we compute the L1 distance between the center coordinates of the predicted bounding box and those of the GT bounding box, providing a direct measure of spatial alignment.
Implementation
We use the AMSGrad optimizer [49, 32] during training and train our models for 30 epochs. Both the transformer encoder and decoder consist of layers, each with attention heads and an MLP dimension of . We apply dropout [54] with in every layer of both the transformer encoder and decoder. The learning rate is set to for the text encoder and for the remaining components. All experiments are conducted on a single NVIDIA RTX A6000 GPU with a batch size of 4.
| CLIP Score | ||||||||
| Objet Sizes | All | S | M | L | All | S | M | L |
| Touch-Line-EWL | 0.7369 | 0.6936 | 0.7305 | 0.7845 | 0.3168 | 0.3006 | 0.2903 | 0.3564 |
| Touch-Line-VTL | 0.7368 | 0.6925 | 0.7321 | 0.7846 | 0.2843 | 0.2809 | 0.2276 | 0.3393 |
| Ours | 0.7983 | 0.7927 | 0.8015 | 0.8010 | 0.2476 | 0.2137 | 0.2241 | 0.3023 |
4.1 Results
Tab. 1 compares the performance of our method with existing approaches on the YouRefIt benchmark. Our model achieves SOTA results across all objects at IoU thresholds of 0.25 and 0.50, despite being trained for only 30 epochs on a single GPU with a batch size of 4. According to our experiments, this limited training setup affects bounding box precision, contributing to the suboptimal performance at , alongside limitations related to the backbone (see Grounding DINO in Tab. 1 for backbone). In contrast, Touch-Line [35] was trained for 200 epochs with a batch size of 56. This clearly demonstrates the efficiency and effectiveness of our approach. Tab. 2 presents a comparison between our method and the previous SOTA, Touch-Line, using two new metrics in this field: CLIP score and . The higher CLIP scores indicate that our predicted bounding boxes have stronger semantic alignment with the input text, suggesting better object identification and improved bounding box precision. Similarly, our method achieves the best performance under the metric, which quantifies the distance between the centers of the predicted and GT bounding boxes. This confirms that our model produces more spatially accurate predictions. In Fig. 3, we present samples images with predicted and GT bounding boxes. We compare our model with Touch-Line [35].
| Setup | Method | IoU=0.25 | IoU=0.5 | IoU=0.75 | CLIP | |
| A | Baseline | 71.2 | 60.1 | 32.8 | 0.7409 | 0.2662 |
| B | A + object center prediction | 70.8 | 61.6 | 34.6 | 0.7339 | 0.2707 |
| C | A + W2F heatmap | 68.9 | 59.7 | 32.4 | 0.7353 | 0.2984 |
| D | A + H2F heatmap | 71.9 | 62.8 | 33.8 | 0.7375 | 0.2694 |
| E | B + W2F heatmap () | 69.6 | 60.7 | 31.5 | 0.7344 | 0.2770 |
| F | B + H2F heatmap () | 72.9 | 62.3 | 35.1 | 0.7373 | 0.2490 |
| G | A + W2F heatmap + H2F heatmap | 70.2 | 60.2 | 33.8 | 0.7345 | 0.2744 |
| H | Full model - Ensemble of E and F | 75.0 | 65.4 | 35.7 | 0.7983 | 0.2476 |
4.2 Analysis
In this section, we provide a detailed analysis of the optimality of our design choices, including heatmap generation and the ensemble method. Specifically, we examine various heatmap injection strategies and ensemble approaches, among other factors. The baseline in Tab. 3 (Setup A) refers to a straightforward training of our model without incorporating any of the proposed contributions.
Object center prediction
The object center prediction head and its associated loss yield improvements over the baseline in terms of mAP at and . However, they have a detrimental effect on mAP at , as well as on the CLIP Score and . Notably, when combined with the heatmap (Setup E and F in Tab. 3), we observe improvements across most metrics (see comparisons from Setup C to E and from Setup D to F).
| IoU=0.25 | IoU=0.5 | IoU=0.75 | CLIP | ||
| Channel-wise input | 68.7 | 58.5 | 33.4 | 0.7348 | 0.2959 |
| Channel-wise feature | 72.6 | 58.4 | 27.9 | 0.7372 | 0.2598 |
| Embedding feature | 72.9 | 62.2 | 35.1 | 0.7373 | 0.2490 |
Where to inject heatmap
We explore various methods for incorporating the heatmap into the model and present the results in Tab. 4. First, we examine the channel-wise input approach, which involves concatenating the heatmap with the input image and feeding a 4-channel input into the image encoder. However, this method leads to a slight drop in performance. This is likely because the image encoder struggles to extract meaningful supervision from the fourth channel, as the heatmap is sparsely represented compared to the image content. In the second approach, channel-wise feature, we adopt the head prompting strategy proposed in [51]. Specifically, we use an embedding layer to learn heatmap embeddings, which are then added to the visual tokens. While this method is computationally efficient and does not require retraining the image encoder, the resulting performance improvement for our task is minimal. Finally, in the embedding feature setup, we introduce an additional CNN-based encoder to process the heatmap. The resulting heatmap embedding is passed through a projection head and then concatenated with the image and text embeddings before being fed into the transformer encoder. This approach yields the most substantial performance improvement among the methods evaluated.
| Ensemble method | IoU=0.25 | IoU=0.5 | IoU=0.75 | CLIP | |
| Confidence-Only | 73.4 | 64.1 | 34.1 | 0.7735 | 0.2470 |
| CLIP-Only (Top-1) | 73.5 | 64.2 | 35.4 | 0.7932 | 0.2391 |
| CLIP-Only (Top-2 + Threshold=0.95) | 73.8 | 64.1 | 35.5 | 0.7938 | 0.2377 |
| CLIP Fusion | 73.6 | 64.4 | 34.5 | 0.7379 | 0.2422 |
| CAPE | 75.0 | 65.4 | 35.7 | 0.7983 | 0.2476 |
CLIP-Aware Pointing Ensemble (CAPE)
We evaluate several ensemble strategies to determine the final referent bounding box from the outputs of the two pointing models and present the results in Tab. 5. (1) Confidence-Only: selects the box with the highest model confidence score among Top-1 prediction of both models. (2) CLIP-Only (Top-1): chooses the box with the highest CLIP similarity score between the text and Top-1 predictions from both models. (3) CLIP-Only (Top-2 + Threshold): extends the previous method by also considering the top-2 predictions from each model, provided their confidence scores exceed . (4) CLIP Fusion: computes a normalized CLIP score and sums it with the model’s confidence score. The box with the highest total score is selected. This method considers the top-2 predictions from both models. (5) CAPE: a hybrid method that combines approaches (3) and (4). For small objects, defined by [11] as occupying less than 0.48% of the image area, CAPE uses method (4), since CLIP tends to perform unreliably on very small referents. For all other objects, it uses method (3). This adaptive strategy balances the strengths of both scoring signals across varying object sizes. Please note that the threshold used in method (3), as well as the decision to apply method (4) for small objects, were determined based on performance analysis on the validation set to avoid overfitting to the test set.
Finetuned vs. frozen text encoder
We investigate the impact of freezing the text encoder during training. While freezing the text encoder reduces computational cost, mitigates overfitting in low-resource settings, and stabilizes the training (see Appendix D), we observe that finetuning consistently yields better performance across almost all evaluation metrics (see Tab. 6). This suggests that adapting the text features to the specific task of embodied reference understanding provides a meaningful benefit.
| Text Encoder | IoU=0.25 | IoU=0.5 | IoU=0.75 | CLIP | |
| Frozen, Setup A | 69.8 | 59.5 | 33.1 | 0.7363 | 0.2994 |
| Finetuned, Setup A | 71.2 | 60.1 | 32.8 | 0.7409 | 0.2662 |
| Frozen, Setup F | 71.3 | 62.1 | 33.7 | 0.7382 | 0.2634 |
| Finetuned, Setup F | 72.9 | 62.2 | 35.1 | 0.7373 | 0.2490 |
How to create heatmap
We evaluate the Conic Attention Heatmap with and cone angles, as well as the Gaussian Ray Heatmap, which yields a more localized activation. Empirically, the Gaussian Ray Heatmap achieved the best performance (see Tab. 7). This suggests that broader attention regions may be less effective for the model, likely due to the high density of objects in typical scenes. Therefore, larger heatmaps tend to enclose multiple distractors, reducing the discriminative value of the spatial signal.
| Heatmap style | IoU=0.25 | IoU=0.5 | IoU=0.75 | CLIP | |
| Gaussian Ray Heatmap () | 72.9 | 62.2 | 35.1 | 0.7373 | 0.2490 |
| Conic Attention Heatmap () | 70.0 | 60.0 | 33.9 | 0.7392 | 0.2632 |
| Conic Attention Heatmap () | 71.5 | 59.4 | 30.8 | 0.7389 | 0.2764 |
5 Conclusion
We tackled the challenge of Embodied Reference Understanding by addressing the limitations of existing single-line pointing assumptions. Our proposed dual-model framework explicitly leverages both head-to-fingertip and wrist-to-fingertip lines, represented as Gaussian ray heatmaps, to provide richer supervision beyond following pointing line and improve pointing target detection. By integrating these complementary signals through our CLIP-Aware Pointing Ensemble (CAPE) module and enhancing the model with an auxiliary object center prediction task, we achieve notable improvements on the YouRefIt benchmark. Our results demonstrate that combining multiple pointing cues and multimodal signals leads to more accurate and robust referent understanding in complex visual scenes.
Limitations
Due to the lack of explicit depth information, the model occasionally selects an incorrect object that lies along the pointing direction but is not the intended referent. We plan to address this depth-related issue in future work. Additionally, the use of a CNN-based image backbone leads to suboptimal precision in bounding box predictions. This limitation can be mitigated by adopting a more advanced backbone architecture, such as the Swin Transformer.
Supplementary Material
Appendix A Problem Analysis
How important is ERU task?
Embodied Reference Understanding (ERU) is a challenging multimodal task that combines visual, linguistic, and embodied cues (like gestures / pointing) to predict a target object in a scene. Despite recent progress, several key problems and open challenges remain. Natural language instructions are often vague or underspecified. Therefore, grounding models may detect multiple candidate objects or fail to resolve the correct referent altogether. Another problem is that when multiple instances of the same object class exist, models struggle to disambiguate. This results in over-detection (selecting all candidates) or under-confidence in selecting the target.
YouRefIt [11] introduced the ERU task, along with a multimodal dataset and a baseline approach to address it. However, the proposed method demonstrated limited performance, achieving a mean Average Precision (mAP) of only under an IoU threshold of . Subsequently, Touch-Line Transformer [35] proposed the concept of a Virtual Touch Line (VTL), based on the observation that the referent is typically aligned with the head-to-fingertip (H2F) line rather than the elbow-to-wrist (E2W) line. Since models struggle to follow pointing instructions, they propose using VTL prediction with a dedicated prediction head to encourage the model to focus more on the corresponding line. This approach led to a significant improvement in performance. However, our empirical analysis confirms that while the referent often aligns with the head-to-fingertip line, this is not always the case due to several factors:
If the person is looking elsewhere, such as at the camera, a robot, or another person, the alignment no longer holds.
The camera angle, consequently the image perspective, affects this alignment.
If the person is very close to the object or does not fully extend their arm while pointing the alignment may again be lost.
As a result, we find that in some cases, the referent aligns with the wrist-to-fingertip line. This indicates the need to incorporate both sources of information either in model training or decision step.
Appendix B Heatmap Generation
The heatmap generation method is a critical component of our approach. Recent work in this field [35] has shown that the referent often aligns with the head-to-fingertip line, referred to as the Virtual Touch-Line (VTL). Based on this insight, we generate a Gaussian Ray Heatmap using the head-to-fingertip line as a reference. While this assumption generally holds true, particularly when a person clearly points toward an object with an extended arm, it is not universally reliable. Variations occur both in this benchmark and in real-world scenarios. For instance, the alignment may be disrupted if the person is looking elsewhere (e.g., at a robot, camera, or another person). Similarly, camera perspective can also distort the alignment. In some cases, we observe that the wrist-to-fingertip line aligns more accurately with the referent than the head-to-fingertip line. To address this variability, we leverage both types of information and generate separate heatmaps for each sample accordingly.
: Heatmap from head-to-fingertip.
: Heatmap from wrist-to-fingertip.
B.1 Head-to-Fingertip Heatmap
To generate this heatmap, we use the head and fingertip as reference points to draw a line starting from the head and extending toward the edge of the image in the direction of the fingertip. In the YouRefIt dataset [11], we have access to detailed annotations from [11, 35], which include coordinates for the eyes, elbows, wrists, and fingertips. We utilize these annotations to construct the reference line and generate the corresponding heatmap. In real-world applications where such annotations are not available, estimating these key points is straightforward. For the eye position, face detection followed by selecting the center point of the detected face yields comparable performance. Likewise, for wrist and fingertip locations, modern pose estimation models provide reliable predictions.
B.2 Wrist-to-Fingertip Heatmap
We follow the same pipeline used for generating the head-to-fingertip heatmap, but replace the eye coordinates with wrist coordinates and the fingertip coordinates remain the same.
B.3 Merging
For the unimodal experiments, we merge the head-to-fingertip and wrist-to-fingertip heatmaps to combine information from different pointing directions. However, this approach yields lower accuracy compared to using two separate models. We hypothesize that this is primarily because the model cannot inherently determine which pointing line is more appropriate for each sample.
B.4 Heatmap Injection Methods
We investigate three different methods for injecting heatmaps and conduct experiments using the head-to-fingertip heatmap.
Channel-wise input
In this method, we concatenate the input image with the heatmap and feed it into the image encoder, resulting in a 4-channel input. Consequently, this approach is not compatible with pretrained image encoders unless they are finetuned, as it requires adapting the model to handle the additional channel. Experimental results show that this method yields the least accurate performance.
Channel-wise feature
In this method, we follow a recently proposed approach [51] that has shown strong performance in the gaze target estimation literature. Specifically, we use a single learnable embedding layer to obtain embeddings of the heatmap, which are then added to the image embeddings.
Embedding feature
In this approach, we propose using an additional encoder. Given the sparse nature of the heatmap representation, we select a simpler encoder than the main image encoder. In this work, we employ ResNet-18 as the heatmap encoder. After obtaining the embedded representation, we pass it through a projection layer to match the required dimensionality. We then concatenate it with the image and text embeddings and feed the combined representation into the transformer encoder.
Appendix C Ensemble Methods
In this section, we present the details of the ensemble methods we explored. Each of our models produces predictions, with each prediction accompanied by a confidence score. Before performing ensembling, we sort these predictions in descending order based on their confidence scores. We denote the sorted prediction list from model as , and from model as . Similarly, we denote the confidence score of model as , and from model as .
C.1 Confidence-Only
In the Confidence-Only approach, we focus solely on the top-1 prediction from each model. We take the bounding box with the highest confidence score from and another from . These two predictions are then compared, and the one with the higher confidence score is selected as the final prediction.
C.2 CLIP-Only (Top-1)
In this method, similar to the Confidence-Only approach, we select the bounding box with the highest confidence score from each model. Next, we compute the CLIP similarity score between each predicted bounding box and the input text. The bounding box with the higher CLIP similarity score is then chosen as the final prediction.
where is the embedded representation of cropped predicted bounding box of with the highest score and is the embedded representation of the text input.
where is the embedded representation of cropped predicted bounding box of with the highest score.
C.3 CLIP-Only (Top-2 + Threshold)
This method is similar to CLIP-Only (Top-1), where we take predictions from both models and compute the CLIP similarity score between the cropped bounding boxes and the input text. However, unlike the Top-1 variant, we also consider the Top-2 predictions from each model. However, we do not always include the second predictions; instead, we apply a confidence threshold determined empirically on the validation set, . If a model’s second-highest prediction has a confidence score greater than or equal to this threshold, it is included in the comparison. Finally, we select the bounding box with the highest CLIP similarity score as the final prediction. The algorithm for this method is presented in Algorithm C.3.
C.4 CLIP Fusion
In this method, we leverage both the CLIP similarity score and the model’s confidence score to make the final prediction. Specifically, we compute a hybrid score for each bounding box by summing its confidence score and its CLIP similarity score. We use the Top-2 predictions from both models for this process. However, the CLIP similarity scores and model confidence scores operate on different scales: while confidence scores are normalized in the range , CLIP similarity scores typically range around . To balance their influence and prevent the CLIP score from dominating, we scale the CLIP similarity scores by a factor of . The scaled CLIP score is then added to the corresponding confidence score to compute the final hybrid score for each bounding box. Finally, we select the bounding box with the highest hybrid score as the final prediction. The algorithm for this method is presented in Algorithm C.4.
C.5 CLIP-Aware Pointing Ensemble (CAPE)
In this method, we combine the CLIP-Only (Top-2 + Threshold) and CLIP Fusion approaches. Through empirical analysis on validation set, we observe that for small objects, the CLIP Fusion method yields more accurate results than other ensemble strategies. This is primarily because CLIP tends to perform less reliably on smaller objects. Therefore, we apply CLIP Fusion to not rely on only CLIP-score but also consider the confidence score. To determine whether an object is small, we follow the definition provided by [11]: if the object occupies less than of the total image area, we classify it as small and use the CLIP Fusion ensemble. For all other cases, we apply the CLIP-Only (Top-2 + Threshold) method. The algorithm is presented in Algorithm C.5.
| Name | Value |
| Learning rate vision backbone | |
| Learning rate - text backbone | |
| Learning rate | 1e-5 |
| 2 | |
| 1 | |
| 10 | |
| 10 | |
| 1 | |
| 1 | |
| Transformer encoder layer | 6 |
| Transformer decoder layer | 6 |
| Number of attention heads per layer | 8 |
| MLP dimension | 2048 |
| Dropout | p = 0.1 |
| Weight decay | 1e-4 |
| Batch size | 4 |
| Text encoder | Roberta-Base |
| Image encoder | ResNet-101 |
| Heatmap encoder | ResNet-18 |
| Position embedding | Sine |
| Input dim of transformer encoder | 256 |
| Number of queries | 20 |
Appendix D Finetuning or Frozen Text Encoder?
Based on multiple training experiments, we observed that finetuning the text encoder sometimes leads to instability, represented by the pink line in Fig. E.5. In contrast, freezing the text encoder results in consistently stable training, as shown by the green and orange lines in Fig. E.5. Therefore, achieving stable training with a finetuned text encoder requires several attempts. However, despite this challenge, finetuning the text encoder ultimately yields better performance.
Appendix E Hyperparameters and Other Training Parameters
In Tab. C.8, we present empirically determined hyperparameters and other key training details to support reproducibility.
Appendix F More Results
In Tab. F.9, we report the mAP at three different thresholds for all ablation setups, categorized by object size. Similarly, in Tab. F.10, we present the CLIP scores and metrics for the same setups, also with respect to object size.
In Fig. F.6, we present examples of failure cases. In the first row, although our model correctly detects the target object, CAPE selects the prediction from , which is incorrect in this case. In the second row, the bottles are extremely difficult to distinguish, even with explicit pointing cues. Nevertheless, the predicted bounding boxes cover other bottles adjacent to the GT one, which is acceptable, as they overlap with the pointing line and are semantically consistent with the text. In the third row, both of our models, as well as the Touch-Line models, produce incorrect predictions. However, guided by the pointing information, the predictions from and lie on the pointing line and are very close to the GT object. In contrast, the Touch-Line predictions are entirely off the pointing line and significantly distant from the GT object.
| IoU Threshold for mAP | 0.25 | 0.50 | 0.75 | ||||||||||||
| Setup | Object Sizes | Heatmap Injection | Ensemble | All | S | M | L | All | S | M | L | All | S | M | L |
| Touch-Line-EWL [35] | - | - | 69.5 | 56.6 | 71.7 | 80.0 | 60.7 | 44.4 | 66.2 | 71.2 | 35.5 | 11.8 | 38.9 | 55.0 | |
| Touch-Line-VTL [35] | - | - | 71.1 | 55.9 | 75.5 | 81.7 | 63.5 | 47.0 | 70.2 | 73.1 | 39.0 | 13.4 | 45.2 | 57.8 | |
| 1 | Baseline | - | - | 71.2 | 59.8 | 73.0 | 80.9 | 60.1 | 43.2 | 66.4 | 70.5 | 32.8 | 8.6 | 35.4 | 53.4 |
| 2 | Setup 1 + object center prediction | - | - | 70.8 | 59.5 | 73.2 | 79.5 | 61.7 | 45.8 | 67.4 | 71.5 | 34.6 | 14.0 | 38.1 | 51.1 |
| 3 | Setup 1 + W2F heatmap | Embedding feature | - | 68.9 | 55.5 | 74.0 | 77.4 | 59.7 | 42.0 | 68.1 | 69.3 | 32.4 | 8.7 | 37.8 | 50.6 |
| 4 | Setup 1 + H2F heatmap | Embedding feature | - | 71.9 | 57.5 | 74.1 | 84.1 | 62.8 | 43.3 | 70.1 | 75.1 | 33.8 | 10.0 | 35.6 | 55.3 |
| 5 | Setup 2 + W2F heatmap | Embedding feature | - | 69.6 | 57.7 | 74.9 | 76.4 | 60.7 | 45.5 | 69.0 | 68.2 | 31.5 | 12.1 | 35.3 | 47.0 |
| 6 | Setup 2 + H2F heatmap | Embedding feature | - | 72.9 | 59.5 | 77.8 | 81.4 | 62.3 | 44.3 | 70.9 | 42.0 | 35.1 | 10.4 | 39.3 | 55.3 |
| 7 | Setup 1 + W2F heatmap + H2F heatmap | Embedding feature | - | 70.2 | 59.0 | 71.4 | 80.2 | 60.2 | 44.5 | 65.6 | 70.5 | 33.8 | 11.9 | 39.3 | 50.0 |
| 8 | Setup 6 | Channel-wise input | - | 68.7 | 58.5 | 71.9 | 75.6 | 58.5 | 44.8 | 65.0 | 65.9 | 33.4 | 13.9 | 37.4 | 48.6 |
| 9 | Setup 6 | Channel-wise feature | - | 72.6 | 60.0 | 77.5 | 80.4 | 58.4 | 38.8 | 68.7 | 68.1 | 27.9 | 8.5 | 32.6 | 42.5 |
| 10 | Setup 5 + Setup 6 | Embedding feature | Confidence-Only | 73.4 | 62.7 | 78.6 | 79.0 | 64.1 | 49.0 | 72.7 | 71.0 | 34.1 | 13.4 | 38.2 | 50.5 |
| 11 | Setup 5 + Setup 6 | Embedding feature | CLIP-Only (Top-1) | 73.5 | 59.7 | 79.9 | 81.2 | 64.2 | 46.5 | 74.3 | 72.2 | 35.4 | 12.7 | 39.8 | 53.4 |
| 12 | Setup 5 + Setup 6 | Embedding feature | CLIP-Only (Top-2 + Threshold) | 73.8 | 59.7 | 80.2 | 81.6 | 64.1 | 46.3 | 74.3 | 72.2 | 35.5 | 12.4 | 40.1 | 53.9 |
| 13 | Setup 5 + Setup 6 | Embedding feature | CLIP Fusion | 73.6 | 63.2 | 78.4 | 79.6 | 64.4 | 49.3 | 72.8 | 71.7 | 34.5 | 13.4 | 38.5 | 51.7 |
| 14 | Setup 5 + Setup 6 | Embedding feature | CAPE | 75.0 | 63.2 | 80.2 | 81.8 | 65.4 | 49.5 | 74.3 | 72.7 | 35.7 | 13.4 | 40.1 | 53.4 |
| 15 | Setup 6 w/ frozen text encoder | Embedding feature | - | 71.3 | 56.0 | 74.9 | 83.1 | 62.1 | 42.5 | 69.5 | 74.4 | 33.7 | 10.2 | 35.8 | 54.6 |
| 16 | Setup 6 w/ Conic Attention Heatmap () | Embedding feature | - | 70.0 | 54.2 | 72.2 | 83.3 | 60.0 | 42.0 | 66.6 | 71.5 | 33.9 | 12.2 | 40.6 | 48.8 |
| 17 | Setup 6 w/ Conic Attention Heatmap () | Embedding feature | - | 71.5 | 58.7 | 73.5 | 82.1 | 59.4 | 41.8 | 64.7 | 71.7 | 30.8 | 7.0 | 32.4 | 52.7 |
| Ours - Final | Embedding feature | CAPE | 75.0 | 63.2 | 80.2 | 81.8 | 65.4 | 49.5 | 74.3 | 72.7 | 35.7 | 13.4 | 40.1 | 53.4 | |
| CLIP Score | |||||||||||
| Setup | Object Sizes | Heatmap Injection | Ensemble | All | S | M | L | All | S | M | L |
| Touch-Line-EWL [35] | - | - | 0.7369 | 0.6936 | 0.7305 | 0.7845 | 0.3168 | 0.3006 | 0.2903 | 0.3564 | |
| Touch-Line-VTL [35] | - | - | 0.7368 | 0.6925 | 0.7321 | 0.7846 | 0.2843 | 0.2809 | 0.2276 | 0.3393 | |
| 1 | Baseline | - | - | 0.7409 | 0.7003 | 0.7341 | 0.7861 | 0.2662 | 0.2182 | 0.2506 | 0.3266 |
| 2 | Setup 1 + object center prediction | - | - | 0.7339 | 0.6947 | 0.7277 | 0.7787 | 0.2707 | 0.2201 | 0.2692 | 0.3222 |
| 3 | Setup 1 + W2F heatmap | Embedding feature | - | 0.7353 | 0.6934 | 0.7297 | 0.7812 | 0.2984 | 0.2763 | 0.2616 | 0.3533 |
| 4 | Setup 1 + H2F heatmap | Embedding feature | - | 0.7375 | 0.6936 | 0.7333 | 0.7852 | 0.2694 | 0.2649 | 0.2233 | 0.3172 |
| 5 | Setup 2 + W2F heatmap | Embedding feature | - | 0.7344 | 0.6939 | 0.7306 | 0.7780 | 0.2770 | 0.2568 | 0.2319 | 0.3382 |
| 6 | Setup 2 + H2F heatmap | Embedding feature | - | 0.7373 | 0.6953 | 0.7343 | 0.7819 | 0.249 | 0.222 | 0.2202 | 0.3024 |
| 7 | Setup 1 + W2F heatmap + H2F heatmap | Embedding feature | - | 0.7345 | 0.6965 | 0.7277 | 0.7780 | 0.2744 | 0.2540 | 0.2314 | 0.3335 |
| 8 | Setup 6 | Channel-wise input | - | 0.7348 | 0.6933 | 0.7273 | 0.7828 | 0.2959 | 0.2437 | 0.2642 | 0.3767 |
| 9 | Setup 6 | Channel-wise feature | - | 0.7372 | 0.6953 | 0.7311 | 0.7839 | 0.2598 | 0.2289 | 0.2119 | 0.3339 |
| 10 | Setup 5 + Setup 6 | Embedding feature | Confidence-Only | 0.7735 | 0.7441 | 0.7843 | 0.7923 | 0.247 | 0.2176 | 0.2042 | 0.3141 |
| 11 | Setup 5 + Setup 6 | Embedding feature | CLIP-Only (Top-1) | 0.7932 | 0.7688 | 0.8010 | 0.8098 | 0.2391 | 0.2242 | 0.2011 | 0.2878 |
| 12 | Setup 5 + Setup 6 | Embedding feature | CLIP-Only (Top-2 + Threshold) | 0.7379 | 0.7967 | 0.7343 | 0.7819 | 0.2422 | 0.2158 | 0.2009 | 0.3060 |
| 13 | Setup 5 + Setup 6 | Embedding feature | CLIP Fusion | 0.7742 | 0.7927 | 0.7455 | 0.7842 | 0.2434 | 0.2137 | 0.2028 | 0.3090 |
| 14 | Setup 5 + Setup 6 (Final Model) | Embedding feature | CAPE | 0.7983 | 0.7927 | 0.8015 | 0.8010 | 0.2476 | 0.2137 | 0.2241 | 0.3023 |
| 15 | Setup 6 w/ frozen text encoder | Embedding feature | - | 0.7382 | 0.6954 | 0.7323 | 0.7855 | 0.2634 | 0.2626 | 0.2233 | 0.3006 |
| 16 | Setup 6 w/ Conic Attention Heatmap () | Embedding feature | - | 0.7387 | 0.6962 | 0.7325 | 0.7860 | 0.2695 | 0.2685 | 0.2344 | 0.3028 |
| 17 | Setup 6 w/ Conic Attention Heatmap () | Embedding feature | - | 0.7391 | 0.6970 | 0.7332 | 0.7863 | 0.2765 | 0.2750 | 0.2499 | 0.3026 |
| Ours - Final | Embedding feature | CAPE | 0.7983 | 0.7927 | 0.8015 | 0.8010 | 0.2476 | 0.2137 | 0.2241 | 0.3023 | |